# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ 6.2 Segmentation Metrics
The common metrics for evaluating segmentation are the Mean Intersection over union (mIoU), pixel accuracy, Average Precision (AP) (refer to section 5.2.1), BF score, and Panoptic Quality. The following sections will explain each, skipping IoU, and AP already discussed.
6.2.1 Pixel Accuracy
@@
Measures the proportion of correctly classified pixels in the whole image. It is calculated by dividing the number of correctly classified pixels by the total number of pixels in the image. The formula for pixel accuracy is
@@
Like regular accuracy, pixel accuracy can be misleading when the class imbalance is high, as it does not consider false positives or negatives. This is why other metrics, such as Intersection over Union (IoU) or Dice coefficient, are commonly used to evaluate image segmentation models.
6.2.2 Boundary F1 Score (BF)
The Boundary F1 Score (BF) [151], often abbreviated as BF score, is a metric used for evaluating image segmentation quality, particularly when the precision of the boundary location is important. The BF score applies the concept of the F1 score to the segmentation boundaries rather than the segment regions themselves.
The computation of the BF score involves the following steps:
@@ 6.2.2 Boundary F1 Score (BF)
    For each segment in the ground truth, find the closest predicted segment (according to some distance measure, often the mean shortest distance between the boundaries of the two segments).
@@
    Compute the precision as the proportion of predicted segments close enough to a ground truth segment. Formally, this is $P=TP/(TP+FP)$, where $TP$ (True Positive) is the number of predicted segments close enough to a ground truth segment, and $FP$ (False Positive) is the number of predicted segments not close enough to any ground truth segment.
@@
    Compute the recall as the proportion of ground truth segments close enough to a predicted segment. Formally, this is $R=TP/(TP+FN)$, where $FN$ (False Negative) is the number of ground truth segments not close enough to any predicted segment.
@@
    The BF score is then the harmonic mean of precision and recall:
$F\_score=\frac{2*Precision*Recall}{Precision+Recall}$ (46)
A threshold is typically defined as close enough when matching predicted segments to ground truth segments. The BF score ranges from 0 to 1, with 1 indicating a perfect match between the predicted and ground truth boundaries.
6.2.3 Panoptic Quality (PQ)
Panoptic Quality (PQ) is a metric proposed for evaluating panoptic segmentation tasks [107].
The Panoptic Quality metric is defined as:
@@ 6.2.3 Panoptic Quality (PQ)
where
  • IoU(p, g) denotes the intersection-over-union of prediction p and ground truth g.
  • TP (True Positive) is a set of matched pairs of predicted and ground truth segments.
  • FP (False Positive) is a set of predicted segments not matched with any ground truth segment.
  • FN (False Negative) is a set of ground truth segments not matched with any predicted segment.
The PQ metric ranges between 0 and 1, where 1 signifies a perfect segmentation. It is a product of two terms:
The first term, called segmentation quality (SQ), calculates the average IoU of the correctly predicted segments (true positives). This term measures how accurately each detected object has been segmented and thus rewards prediction quality.
The second term, called Recognition Quality (RQ), calculates the ratio of the number of true positives to the total number of segments. This term measures how accurately each object has been recognized.
This metric can be more informative than mean IoU when dealing with complex scenes containing multiple object instances per class, as it considers both the segmentation quality and the correct recognition of distinct instances.
7 Face Recognition
Face recognition aims to accurately match an individual’s face in an image or video to a corresponding entry in a database of faces. This task is often performed using deep learning algorithms, such as Convolutional Neural Networks (CNNs) or Transformers [152], that are trained on large datasets of face images. The algorithms are trained to extract features from images faces and then use these features to recognize faces that match those in the database. Face recognition has many applications, including security [153, 154], social media [155], and biometric identification systems [156, 157].
7.1 Face Recognition Loss Functions and Metrics
@@
7.1.1 Softmax Loss
The Softmax Loss computes the cross-entropy between the predicted class probabilities and the true class label, followed by the logarithmic operation to convert the negative log-likelihood into a loss value. The final loss is obtained by summing the cross-entropy over all classes and all samples.
Let’s denote the weight vectors for each class (or, in this case, each face identity) as $W=\{w_{1},w_{2},...,w_{n}\}$, where $n$ is the number of classes (or identities). For a given input image $x$ of class $y$, the linear classifier would compute a score $f(x,W)=Wx+b$, where $b$ is the bias term.
The softmax function converts these scores into probabilities. The probability of the $i^{th}$ class is computed as follows:
@@ 7.1.1 Softmax Loss
The softmax loss (also known as the cross-entropy loss) for an input-output pair $(x,y)$ is defined as the negative log-likelihood of the correct class, which we can be expressed as
@@
where $f_{y_{i}}$ is the score for the correct class, and $f_{j}$ are the scores for all classes. The total loss for a batch of data is the mean of $L_{i}$ over all the examples in the batch.
The disadvantage of this loss function is that it does not have fine-grained control over the intra-class and inter-class distances that come in handy for face recognition purposes.
7.1.2 A-Softmax Loss
The A-Softmax loss [158], also known as the SphereFace loss, was designed to address the limitations of the traditional softmax loss by considering the angular information between the features of face images and their corresponding labels. The A-Softmax loss aims to maximize the inter-class separability and minimize the intra-class variations of face features.
@@ 7.1.2 A-Softmax Loss
1. Compute the Normalized weight matrix $W_{norm}$ as
$W_{norm}=\frac{W}{||W||}$ (50)
where each column $w_{i}$ in $W_{norm}$ is a unit vector, i.e., $||w_{i}||=1$. The normalization operation makes the weights for each class lie on the unit hypersphere.
2. Compute the margin $M$ to be applied to the angles:
$M=(m-1)\cdot y_{true}+1$ (51)
In this equation, $y_{true}$ is the true class label. If $y_{true}$ equals 1, $M$ equals $m$, and if $y_{true}$ equals 0, $M$ equals 1.
3. Compute the cosine of the angle $\theta$ between the feature vector $x$ and the weight vector:
$cos(\theta)=\frac{W_{norm}\cdot x}{||x||},$ (52)
@@ 3. Compute the cosine of the angle $\theta$ between the feature vector $x$ and the weight vector:
4. Compute the new angle $\theta^{\prime}$ and its cosine after applying the margin:
$\theta^{\prime}=\theta\cdot M$ (53)
$cos(\theta^{\prime})=cos(\theta^{\prime})$
5. Compute the new prediction $y_{pred}^{\prime}$ by rescaling with the norm of $x$:
$y_{pred}^{\prime}=||x||\cdot cos(\theta^{\prime})$ (54)
6. Finally, compute the SphereFace loss $L$, which is the negative log-likelihood of the true class:
$L=-\log\frac{\sum y_{true}\cdot e^{y_{pred}^{\prime}}}{\sum e^{y_{pred}^{\prime}}}$ (55)
@@ 6. Finally, compute the SphereFace loss $L$, which is the negative log-likelihood of the true class:
Here, the summation is taken over all classes. The numerator is the exponentiated prediction for the true class, and the denominator is the sum of exponentiated predictions for all classes.
Compared to the traditional softmax loss, the A-Softmax loss produces more discriminative and compact features, improving face recognition performance.
7.1.3 Center Loss
The center loss [159] aims to increase the inter-class variance while reducing the intra-class variance by penalizing the distances between the features of a sample and the center of its corresponding class. The center loss is inspired by the idea of having a center for each class in the feature space, and the loss function encourages the features of the same class to be close to its center. The center loss is defined as the Euclidean distance between the feature of a sample and the center of its class in the feature space.
The center loss is usually added to the main loss function in the training process, and it is defined as
$L_{center}=\frac{1}{2}\sum_{i=1}^{n}\left|\mathbf{x_{i}}-\mathbf{c_{y_{i}}}\right|_{2}^{2},$ (56)
where $\mathbf{x_{i}}$ is the feature representation of the $i^{th}$ sample, $y_{i}$ is its corresponding class label, $\mathbf{c_{y_{i}}}$ is the center of class $y_{i}$, and $n$ is the number of samples.
7.1.4 CosFace: Large-Margin Cosine Loss
CosFace loss, also known as the Large Margin Cosine Loss, maximizes the decision margin in the cosine space to further enhance the discriminative power of the deeply learned features.
The cosine of the angle between the feature vector $x_{i}$ and the weight $w_{j}$ of the $j^{th}$ class is given by
$cos\theta_{j}=\frac{w_{j}^{T}x_{i}}{||w_{j}||||x_{i}||},$ (57)
where ||.||||.||| | . | | denotes the l2 norm.
The CosFace method adds a cosine margin $m$ to the target logit, so the modified cosine of the angle $\theta_{y_{i}}$ corresponding to the ground truth class $y_{i}$ is given by:
$cos\theta_{y_{i}}-m$ (58)
Then the CosFace loss for an input-output pair $(x_{i},y_{i})$ is defined as
$L_{i}=-\log\frac{e^{s(cos\theta_{y_{i}})-m}}{e^{s(cos\theta_{y_{i}})-m}+\sum_{j\neq y_{i}}^{n}e^{scos\theta_{j}}},$ (59)
where $s$ is a scaling factor.
One advantage of this function is that the cosine function’s non-monotonicity does not create a problem here, unlike SphereFace. Also, because the feature vector is normalized, the model must learn better separation of the angles as it cannot reduce loss by learning a different norm [160].
7.1.5 ArcFace. Additive Angular Margin Loss
The ArcFace loss [161] enhances the discriminative power of the softmax loss by adding an angular margin penalty to the target logit.
The ArcFace method adds an additive angular margin $m$ to the target logit, so the modified cosine of the angle $\theta_{y_{i}}$ corresponding to the ground truth class $y_{i}$ is given by
$cos(\theta_{y_{i}}+m)$ (60)
Then the ArcFace loss for an input-output pair $(x_{i},y_{i})$ is defined as
$L_{i}=-\log\frac{e^{scos(\theta_{y_{i}}+m)}}{e^{scos(\theta_{y_{i}}+m)}+\sum_{j\neq y_{i}}^{n}e^{scos\theta_{j}}},$ (61)
where $s$ is a scaling factor. The margin $m$ can be interpreted as an additional arc length on the hypersphere of radius $s$. Experiments show better inter-class discrepancy than Triplet Loss while having about the same intra-class similarity [160].
7.1.6 Triplet Loss
Probably the best-known loss function for face recognition. The idea behind the triplet loss [162, 163] is to train the model to distinguish between a positive pair of images (two images of the same person) and a negative pair of images (two images of different persons).
Given an anchor image, $A$, a positive image, $P$, and a negative image, $N$, (see Figure 5), the loss is calculated as the distance between the anchor image and positive image and the distance between the anchor image and negative image, plus a margin. The equation is defined as
$L_{triplet}=\max{0,\left\lVert\textbf{f}_{A}-\textbf{f}_{P}\right\rVert^{2}_{2}-\left\lVert\textbf{f}_{A}-\textbf{f}_{N}\right\rVert^{2}_{2}+\alpha},$ (62)
where $\textbf{f}_{A}$, $\textbf{f}_{P}$, and $\textbf{f}_{N}$ are the embeddings of the anchor image, positive image, and negative image, respectively, and $\alpha$ is a hyperparameter known as the margin. By minimizing this loss, the embeddings of the positive images get closer to each other than the embeddings of the negative images.
Figure 5: Triplet Loss. Given an anchor image, a positive image, and a negative image, the triplet loss computes the distance between the embeddings of the anchor image and positive image and the distance between the embeddings of the anchor image and negative image. By minimizing this loss, the embeddings of the positive images get closer to each other than the embeddings of the negative images.
7.1.7 Contrastive Loss
The contrastive loss [164] learns a feature representation or image embedding that projects similar images close to each other and dissimilar images far apart. This loss is based on a Siamese architecture of the neural network. The contrastive loss function is defined as:
$L=\frac{1}{N}\sum_{i=1}^{N}y_{i}\left|f(x_{i}^{a})-f(x_{i}^{p})\right|_{2}^{2}+(1-y_{i})\max\left(0,m-\left|f(x_{i}^{a})-f(x_{i}^{n})\right|_{2}^{2}\right),$ (63)
where $f(x_{i}^{a})$ is the deep feature representation of the anchor image $x_{i}^{a}$, $f(x_{i}^{p})$ is the deep feature representation of the positive image $x_{i}^{p}$, $f(x_{i}^{n})$ is the deep feature representation of the negative image $x_{i}^{n}$. $y_{i}$ is a binary label indicating whether the anchor and positive images are similar (1) or dissimilar (0), $m$ is a margin hyperparameter, and $N$ is the number of triplets in the batch.
The margin $m$ controls how hard the model should work to push dissimilar embeddings apart. Extending a trained model for new/unseen classes is easy because it learns to create a semantic representation of the image rather than classify it among a predetermined set of classes.
One limitation of this loss is that the margin $m$ is the same constant for all dissimilar pairs, which implicitly pushes the model to have the same distance between all dissimilar pairs, even if some are more dissimilar [160, 165]. A second limitation is that the absolute notion of similar and dissimilar pairs is not transferable from one context to another [166].
7.1.8 Circle Loss
The Circle Loss [167] pushes positive pairs closer and negative pairs farther away while maintaining a circle-like decision boundary. This is achieved by adding margins to positive and negative pairs, enlarging the intra-class variance for negative pairs, and reducing the intra-class variance for positive pairs. By doing so, Circle Loss can effectively handle imbalanced data and complex distributions, which are common challenges in face recognition tasks. The circle loss equation can be expressed as
$\alpha_{pos_{i}}$ $=\max(O_{pos_{i}}-s_{pos_{i}},0)$ (64)
$\alpha_{neg_{j}}$ $=\max(O_{neg_{j}}-s_{neg_{j}},0)$
$sum_{pos}$ $=\sum_{i}e^{-\gamma\cdot\alpha_{pos_{i}}\cdot_{pos_{i}}}$
$sum_{neg}$ $=\sum_{j}e^{\gamma\cdot\alpha_{neg_{j}}\cdot s_{neg_{j}}}$
$L$ $=\log(1+sum_{pos}\cdot sum_{neg})$

### v3

@@ 4.3.2 Segmentation Metrics
The common metrics for evaluating segmentation are the Mean Intersection over union (mIoU) (section 4.3.1), pixel accuracy, Average Precision (AP) (section 4.2.2), BF score, and Panoptic Quality. The following sections will explain each, skipping IoU, and AP already discussed.
Table 13: Common metrics used in different types of segmentation tasks.
Metric Name Type of
Segmentation
Description
Pixel Accuracy Semantic
Segmentation
Measures the percentage of correctly labeled pixels. Best used when class distributions are balanced.
Intersection over Union (IoU) Semantic
Segmentation
Evaluates the overlap between the predicted segmentation and the ground truth over their union. Useful across various classes, particularly when dealing with unbalanced datasets.
Boundary F1 Score Semantic
Segmentation
Measures the harmonic mean of precision and recall for boundary prediction accuracy, particularly useful for assessing the delineation of object boundaries.
Dice Coefficient Semantic
Segmentation
Similar to IoU but considers twice the intersection over the sum of the sizes of both the prediction and the ground truth. Effective for unbalanced datasets.
Precision and Recall Instance
Segmentation
Measure the accuracy and completeness of the detected instances, respectively. Employed to evaluate the detection of instances at various IoU thresholds.
Average Precision (AP) Instance
Segmentation
Computes the average precision across different IoU thresholds; used in competitions and benchmarks like COCO [119].
Mask Mean Average Precision (Mask mAP) Instance Segmentation Assesses the quality of the instance masks generated by the model. It averages the AP across all classes and IoU thresholds.
Panoptic Quality (PQ) Panoptic
Segmentation
Combines recognition and segmentation accuracy into a single metric by multiplying segmentation quality (SQ) by recognition quality (RQ).
Segmentation Quality (SQ) Panoptic
Segmentation
Measures the IoU of correctly matched segments to their corresponding ground truth objects, reflecting the segmentation accuracy.
Recognition Quality (RQ) Panoptic
Segmentation
Evaluates how effectively the model identifies and differentiates instances, independent of segmentation performance.
Pixel Accuracy
@@
Pixel Accuracy measures the proportion of correctly classified pixels in the entire image. It is calculated by dividing the number of correctly classified pixels by the total number of pixels in the image. The formula for Pixel Accuracy is:
@@
While Pixel Accuracy provides a straightforward assessment of overall model performance, its reliability as a sole metric is compromised in cases of high class imbalance. This metric tends to favor the majority class, potentially overlooking misclassified instances of less frequent classes. Therefore, it is often complemented by other metrics such as Intersection over Union (IoU) and the Dice coefficient, which provide insights into the model’s performance across different classes and are less sensitive to class imbalance. Pixel Accuracy might still be useful in scenarios where classes are fairly balanced, and the primary interest is in the general accuracy of pixel classification across the image.
Boundary F1 Score (BF)
The Boundary F1 Score (BF) [178], commonly known as the BF score, evaluates the precision of image segmentation, specifically focusing on the precision of boundary delineation rather than the segmentation of regions. Unlike metrics such as Intersection over Union (IoU), which assess the overlap between predicted and ground truth regions, the BF score is tailored to applications where accurate boundary detection is crucial, such as in medical image analysis.
The calculation of the BF score involves the following steps:
@@
    Identify the closest predicted segment for each ground truth segment using a distance measure. This typically involves calculating the mean shortest distance between each point on the predicted boundary and the closest point on the ground truth boundary, as well as vice versa.
@@
    Calculate the precision as the proportion of predicted segments that are within a predefined distance threshold from any ground truth segment. This is given by $P=\frac{TP}{TP+FP}$, where $TP$ (True Positives) represents the count of predicted segments adequately close to the ground truth, and $FP$ (False Positives) denotes the predicted segments that fail to meet this criterion.
@@
    Determine the recall as the proportion of ground truth segments that are adequately approximated by any predicted segment. This is given by $R=\frac{TP}{TP+FN}$, where $FN$ (False Negatives) refers to ground truth segments that are not adequately approximated by any predicted segment.
@@
    The BF score is then computed as the harmonic mean of precision and recall:
$F_{s}core=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$ (67)
The distance threshold, which defines what is considered close enough, is typically determined based on a pixel distance or another spatial metric, depending on the application and image resolution. The BF score ranges from 0 to 1, with 1 indicating perfect congruence between the predicted and actual boundaries. It is particularly valuable in applications where boundary precision is critical. However, it is important to note that the choice of threshold can significantly impact the score, and the BF score may not fully reflect the quality of region segmentation.
Panoptic Quality (PQ)
Panoptic Quality (PQ) is a comprehensive metric proposed for evaluating panoptic segmentation tasks, which involve both semantic and instance segmentation [134]. This metric accounts for the quality of both the segmentation and the recognition of individual instances.
The PQ metric is defined as:
@@
where:
  • IoU(p, g) denotes the intersection-over-union between a predicted segment p and the corresponding ground truth segment g.
  • TP (True Positives) is the set of matched pairs of predicted and ground truth segments, where a match is determined based on a minimum IoU threshold (commonly 0.5).
  • FP (False Positives) is the set of predicted segments that do not match any ground truth segment.
  • FN (False Negatives) is the set of ground truth segments that are not matched by any predicted segment.
The PQ metric ranges from 0 to 1, with 1 indicating perfect segmentation and recognition. It is the product of two components:
  • Segmentation Quality (SQ): This component measures the average IoU of the correctly detected segments (true positives). It reflects how accurately each correctly predicted object has been segmented, emphasizing the quality of the segment boundaries and overlap.
  • Recognition Quality (RQ): This component measures the ratio of true positives to the sum of true positives, false positives, and false negatives. It captures the accuracy in recognizing and matching the instances of objects, thus balancing precision and recall in detection.
The PQ metric provides a more integrated evaluation compared to the mean IoU, especially in complex scenes with multiple object instances per class. By considering both segmentation accuracy and instance recognition, PQ ensures that models are evaluated not only on the precision of segment boundaries but also on the ability to correctly distinguish and identify all instances within a scene. This makes it particularly valuable in real-world applications where both aspects are critical.
4.4 Face Recognition
Face recognition is the process of accurately matching an individual’s face in an image or video to a corresponding entry in a database of faces. This is often done using deep learning algorithms, such as Convolutional Neural Networks (CNNs) or Transformers [179], that have been trained on large datasets of face images. These algorithms are trained to extract features from images of faces, and then use these features to recognize faces that match those in the database. Face recognition has many applications, including security [180, 181], social media [182], and biometric identification systems [183, 184].
4.4.1 Face Recognition Loss Functions and Metrics
@@
Softmax Loss
Softmax Loss computes the cross-entropy between the predicted class probabilities and the true class labels, transforming the negative log-likelihood into a loss value. The final loss is obtained by summing the cross-entropy over all classes and samples.
Let the weight vectors for each class (or each face identity) be denoted as $W=\{w_{1},w_{2},...,w_{n}\}$, where $n$ is the number of classes (or identities). For a given input image $x$ belonging to class $y$, the linear classifier computes a score $f(x,W)=Wx+b$, where $b$ is the bias term. Here, $w_{i}^{T}x$ represents the dot product between the weight vector $w_{i}$ and the feature vector $x$.
The softmax function converts these scores into probabilities. The probability of the $i^{th}$ class is computed as:
@@
The softmax loss (also known as cross-entropy loss) for an input-output pair $(x,y)$ is defined as the negative log-likelihood of the correct class:
@@
where $f_{y_{i}}$ is the score corresponding to the true class $y_{i}$, and $f_{j}$ are the scores for all classes. The total loss for a batch of data is the mean of $L_{i}$ over all examples in the batch.
However, Softmax Loss has limitations in face recognition tasks as it does not explicitly enforce fine-grained control over intra-class and inter-class distances. This can hinder its effectiveness in distinguishing between similar faces and ensuring that different identities are well-separated in the feature space.
A-Softmax Loss
The A-Softmax loss [185], also known as the SphereFace loss, is designed to address the limitations of traditional softmax loss by incorporating angular information between the feature vectors of face images and their corresponding weight vectors. The A-Softmax loss enhances inter-class separability and reduces intra-class variations by enforcing angular margins, leading to more discriminative features.
@@
1. Compute the normalized weight matrix $W_{norm}$:
$W_{norm}=\frac{W}{||W||},$ (71)
where each column vector $w_{i}$ in $W_{norm}$ is a unit vector, i.e., $||w_{i}||=1$. This normalization ensures that each class’s weight vector lies on the unit hypersphere.
2. Compute the cosine of the angle $\theta$ between the feature vector $x$ and the normalized weight vector:
$\cos(\theta)=\frac{W_{norm}\cdot x}{||x||},$ (72)
@@ 1. Compute the normalized weight matrix $W_{norm}$:
3. Apply the angular margin $m$ to the angle $\theta$. The angle margin enforces that the decision boundary becomes more stringent, requiring the correct class to have a larger cosine similarity:
$\cos(m\theta)=\cos(m\cdot\arccos(\cos(\theta))),$ (73)
where $\arccos(\cos(\theta))$ computes the angle $\theta$ from the cosine value.
4. Compute the scaled prediction logits:
$y_{pred}^{\prime}=||x||\cdot\cos(m\theta).$ (74)
5. The final A-Softmax loss $L$ is calculated as the negative log-likelihood of the correct class:
$L=-\log\frac{e^{y_{pred}^{\prime}}}{e^{y_{pred}^{\prime}}+\sum_{j\neq y}e^{||x||\cdot\cos(\theta_{j})}},$ (75)
@@ 5. The final A-Softmax loss $L$ is calculated as the negative log-likelihood of the correct class:
where the numerator represents the exponentiated scaled prediction for the true class, and the denominator sums the exponentiated scaled predictions for all classes.
The A-Softmax loss introduces an angular margin between different classes, making the features more discriminative. It helps to produce more compact intra-class representations and maximizes inter-class separability, which is particularly beneficial for face recognition tasks, as it enhances the model’s ability to differentiate between different identities.
Center Loss
The center loss [186] aims to reduce intra-class variance while increasing inter-class variance in feature space by minimizing the distances between feature representations of samples and their respective class centers. This loss function is designed to encourage samples from the same class to cluster closely around a central point, known as the class center, thus improving the discriminative power of the learned features.
The center loss is defined as the Euclidean distance between the feature representation of a sample and the corresponding class center in the feature space. It is typically combined with the primary loss function, such as softmax loss, to ensure both discriminative feature learning and correct classification.
The mathematical formulation of the center loss is:
$L_{center}=\frac{1}{2}\sum_{i=1}^{n}\left\|\mathbf{x_{i}}-\mathbf{c_{y_{i}}}\right\|_{2}^{2},$ (76)
@@
  • $\mathbf{x_{i}}$ is the feature representation of the $i^{th}$ sample.
  • $y_{i}$ is the corresponding class label for the $i^{th}$ sample.
  • $\mathbf{c_{y_{i}}}$ is the center of class $y_{i}$ in the feature space.
  • $n$ is the total number of samples in the batch.
The class centers $\mathbf{c_{y_{i}}}$ are updated during the training process. They are calculated by taking the mean of the feature representations of all samples belonging to the same class in the current mini-batch. To prevent abrupt updates and ensure stability, a learning rate parameter $\alpha$ is often used to update the centers incrementally:
$\mathbf{c_{y_{i}}}^{(t+1)}=\mathbf{c_{y_{i}}}^{(t)}+\alpha\left(\mathbf{x_{i}}-\mathbf{c_{y_{i}}}^{(t)}\right),$ (77)
where $\mathbf{c_{y_{i}}}^{(t)}$ is the class center at the $t^{th}$ iteration, and $\alpha$ is the learning rate for the center updates.
The total loss function for the training process is typically a combination of the primary loss (e.g., softmax loss) and the center loss:
$L=L_{softmax}+\lambda L_{center},$ (78)
where $\lambda$ is a weight parameter that balances the influence of the center loss against the primary loss.
The inclusion of center loss helps the model produce features that are not only well-separated across different classes but also tightly clustered within the same class. This is particularly beneficial for tasks like face recognition, where distinguishing between individuals is critical.
CosFace: Large-Margin Cosine Loss
CosFace loss [187], also known as the Large Margin Cosine Loss, is designed to enhance the discriminative power of deeply learned features by maximizing the decision margin in the cosine space.
The cosine of the angle between the feature vector $\mathbf{x_{i}}$ and the weight vector $\mathbf{w_{j}}$ of the $j^{th}$ class is computed as:
$\cos\theta_{j}=\frac{\mathbf{w_{j}}^{T}\mathbf{x_{i}}}{||\mathbf{w_{j}}||\,||\mathbf{x_{i}}||},$ (79)
where |||||$\cdot$||| | ⋅ | | denotes the L2 norm.
To introduce a margin in the cosine space, CosFace modifies the cosine similarity corresponding to the ground truth class $y_{i}$ by subtracting a cosine margin $m$. The adjusted cosine similarity for the ground truth class is given by:
$\cos\theta_{y_{i}}-m.$ (80)
The CosFace loss for an input-output pair $(\mathbf{x_{i}},y_{i})$ is then defined as:
$L_{i}=-\log\frac{e^{s(\cos\theta_{y_{i}}-m)}}{e^{s(\cos\theta_{y_{i}}-m)}+\sum_{j\neq y_{i}}^{n}e^{s\cos\theta_{j}}},$ (81)
where $s$ is a scaling factor that helps in optimizing the convergence of the training process by adjusting the distribution of logits.
The key advantage of CosFace is its ability to enforce a large margin between classes in the normalized cosine space, which leads to better separation of features across different classes. Unlike SphereFace, which uses angular margins, CosFace directly manipulates the cosine similarity, avoiding potential issues related to the non-monotonicity of the cosine function. Additionally, since both the feature vectors and the weight vectors are normalized, the model cannot reduce the loss by simply scaling the norms, thereby ensuring that discriminative learning focuses on the angular separation of features.
ArcFace: Additive Angular Margin Loss
The ArcFace loss [188], also known as Additive Angular Margin Loss, enhances the discriminative power of the softmax loss by incorporating an angular margin penalty into the target logit. This technique aims to achieve better class separation by adding an additional margin in the angular space.
The ArcFace method introduces an additive angular margin $m$ to the angle $\theta_{y_{i}}$ corresponding to the ground truth class $y_{i}$. The modified cosine similarity for the ground truth class is defined as:
$\cos(\theta_{y_{i}}+m),$ (82)
where $\theta_{y_{i}}$ is the angle between the feature vector $\mathbf{x_{i}}$ and the weight vector $\mathbf{w_{y_{i}}}$ corresponding to the true class $y_{i}$.
The ArcFace loss for an input-output pair $(\mathbf{x_{i}},y_{i})$ is then defined as:
$L_{i}=-\log\frac{e^{s\cos(\theta_{y_{i}}+m)}}{e^{s\cos(\theta_{y_{i}}+m)}+\sum_{j\neq y_{i}}^{n}e^{s\cos\theta_{j}}},$ (83)
where $s$ is a scaling factor that controls the smoothness of the decision boundary.
The introduction of the margin $m$ effectively increases the decision margin in the angular space, pushing the features of different classes further apart while keeping features of the same class closely grouped. This margin can be interpreted as an additional arc length on the hypersphere, which the feature vectors and weight vectors are projected onto, with a radius determined by the scaling factor $s$. The angular nature of the margin ensures that the decision boundaries are determined purely by the angles, making the feature separation more robust to variations in feature magnitude.
ArcFace has been shown to provide superior performance in face recognition tasks compared to other loss functions like Triplet Loss, achieving better inter-class discrepancy (separation between different classes) while maintaining high intra-class similarity (compactness within the same class) [189]. This characteristic makes ArcFace particularly effective for applications that require high precision to distinguish between closely related classes, such as identity verification and authentication systems.
Triplet Loss
Triplet Loss is one of the most well-known loss functions for face recognition, designed to encourage the model to distinguish between images of the same person and images of different persons. The core idea behind Triplet Loss [190, 191] is to ensure that the distance between an anchor image and a positive image (of the same person) is smaller than the distance between the anchor image and a negative image (of a different person) by at least a predefined margin.
Given an anchor image, $A$, a positive image, $P$, and a negative image, $N$, the loss function is defined as
$L_{triplet}=\max\{0,\left\lVert\mathbf{f}_{A}-\mathbf{f}_{P}\right\rVert^{2}_{2}-\left\lVert\mathbf{f}_{A}-\mathbf{f}_{N}\right\rVert^{2}_{2}+\alpha\},$ (84)
where $\mathbf{f}_{A}$, $\mathbf{f}_{P}$, and $\mathbf{f}_{N}$ are the feature embeddings of the anchor, positive, and negative images, respectively, and $\alpha$ is the margin parameter. This margin ensures that the model maintains a certain minimum distance between the embeddings of different classes.
The term $\left\lVert\mathbf{f}_{A}-\mathbf{f}_{P}\right\rVert^{2}_{2}$ represents the squared L2 distance between the anchor embeddings and the positive image, while $\left\lVert\mathbf{f}_{A}-\mathbf{f}_{N}\right\rVert^{2}_{2}$ represents the squared L2 distance between the embeddings of the anchor and the negative image.
The goal of the triplet loss is to minimize this loss function, which effectively pulls the positive pairs closer together in the embedding space and pushes the negative pairs farther apart. By doing so, the model learns a feature space where images of the same person have similar embeddings, and images of different persons have distinctly separate embeddings.
The choice of the margin $\alpha$ is crucial; if it is too small, the model might not learn enough discriminative features, whereas if it is too large, it can lead to slow convergence or even instability during training. Careful selection and tuning of $\alpha$ are necessary to achieve optimal performance.
Contrastive Loss
The contrastive loss [192] is designed to learn a feature representation or embedding that projects similar images close to each other and dissimilar images far apart in the feature space. This loss is commonly used with a Siamese neural network architecture, which processes pairs of images to learn these relationships.
The contrastive loss function is defined as
$L=\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\left\|f(x_{i}^{a})-f(x_{i}^{p})\right\|_{2}^{2}+(1-y_{i})\max\left(0,m-\left\|f(x_{i}^{a})-f(x_{i}^{n})\right\|_{2}^{2}\right)\right],$ (85)
where:
  • $f(x_{i}^{a})$ is the feature representation of the anchor image $x_{i}^{a}$.
  • $f(x_{i}^{p})$ is the feature representation of the positive image $x_{i}^{p}$.
  • $f(x_{i}^{n})$ is the feature representation of the negative image $x_{i}^{n}$.
  • $y_{i}$ is a binary label indicating whether the pair (anchor and positive) is similar (1) or dissimilar (0).
  • $m$ is the margin hyperparameter that determines the minimum distance required between dissimilar pairs.
  • $N$ is the number of image pairs in the batch.
The term $\left\|f(x_{i}^{a})-f(x_{i}^{p})\right\|_{2}^{2}$ measures the squared Euclidean distance between the feature representations of the anchor and positive images. For similar pairs, the goal is to minimize this distance. In contrast, the term $\max\left(0,m-\left\|f(x_{i}^{a})-f(x_{i}^{n})\right\|_{2}^{2}\right)$ ensures that the distance between two pairs is at least the margin $m$. If the distance is already greater than $m$, the contribution to the loss is zero.
The margin $m$ acts as a threshold and controls how aggressively the model should separate dissimilar embeddings. The choice of $m$ affects the model’s ability to differentiate between similar and dissimilar pairs. One advantage of contrastive loss is that it facilitates extending a trained model to new or unseen classes, as it focuses on learning a meaningful feature space rather than classifying it into predefined categories.
One limitation of this loss function is that the margin $m$ is a constant for all dissimilar pairs, which may not account for varying degrees of dissimilarity; the model might implicitly assume equal dissimilarity between all negative pairs. Additionally, the absolute notion of similarity and dissimilarity defined by the label $y_{i}$ may not generalize well across different contexts, as what is considered similar in one context may not be perceived the same in another [189, 193, 194].
Circle Loss
Circle loss [195] enhances the discriminative power of learned features by simultaneously maximizing the similarity of positive pairs and minimizing the similarity of negative pairs, while maintaining a circle-like decision boundary. This approach addresses challenges such as imbalanced data and complex distributions, which are prevalent in tasks like face recognition.
The Circle loss is defined as:
$\alpha_{pos_{i}}$ $=\max(O_{pos}-s_{pos_{i}},0)$ (86)
$\alpha_{neg_{j}}$ $=\max(s_{neg_{j}}-O_{neg},0)$
$\text{sum}_{pos}$ $=\sum_{i}e^{-\gamma\cdot\alpha_{pos_{i}}\cdot(s_{pos_{i}}-O_{pos})}$
$\text{sum}_{neg}$ $=\sum_{j}e^{\gamma\cdot\alpha_{neg_{j}}\cdot(s_{neg_{j}}-O_{neg})}$
$L$ $=\log(1+\text{sum}_{pos}\cdot\text{sum}_{neg})$
Where: